OpenStudy (anonymous):

HELP WITH ARITHMETIC SEQUENCES: Andrea is designing the seating arrangement for a concert in her local park. To give everyone a good view, each row must have 4 more seats than the row before it, and the 1st row can only have 10 seats. Help Andrea plan the rest of the seating by solving for how many seats are in row 20. Then explain to Andrea how to create an equation to predict the number of seats in any row. Show your work, and use complete sentences.

OpenStudy (anonymous):

@SolomonZelman @Data_LG2 @Compassionate @batman19991 @Keigh2015 @uri @Elsa213 @iihoop3 @iiAriana @Empty @wil476003 @WazapxD @QueenSid @tatiachrikishvili @Goku-Kai @gabsss @favoreke @theresa2314

OpenStudy (keigh2015):

I don't know I am sorry

OpenStudy (batman19991):

4+4+4+4+4+4+.....n times = 4n. Subtract the first case 4n-4 = 4(n-1) i think

OpenStudy (anonymous):

so we know that S(1) = 10. now we know that S( i ) = S(i-1) + 4. this solves to S(i) = 10 + 4 (i-1). this is true since when i = 1, S = 10, and after that you add 4 each time. so S(20) = 10 + 4(19) = 86.

OpenStudy (anonymous):

@alexistheking777

OpenStudy (anonymous):

The answer is 86 because S(i) =S(i-1)+4 and because S is 10 you and 4 each time so basically S(20)=10+4(19) which equals 86